I have to find the domain for the function. I have no clue on how to do this so any help will be greatly apreciated. f(x)=(x^2-x-12)/(x-2)
Answer by oscargut(2103)

The domain for a function (real) are all values of x (real) where exists the function
her e you have a rational expression so in this case you have to find the zeros of the denominator (i can not divide by 0) so
if f(x)=(x^2-x-12)/(x-2) doing x-2=0 then x=2 ( 2 does not belong to the domain) so Domain is all the real numbers without 2